Solve the following differential equations.
10. Solve the following differential equations. (a) (x2 - y2) 2 = ry (c) y – y cot = cot x (d) - 2y = 23
Answer #1

ng @ Let (x392) dec =xy face) D. (x²y) dy _aydx=0 This is of the form Mdet rioy =0 M --14 N-a?y? Do By 5°N condition For exadd dy-y=elyt - @ This is Bernaullis eq? nad put z=4 -4 = your man = -344 dy z dy pod kde ga3 .:65 = zbytdz -Yety! P e Multiple - e - e 7 dye-?y=22 - This is linear diffeq? - 1-2 dx -zloga I-f=ě a solution is Jaya Step-d8e + c = la dato I try = azt 1